Operations Report – December 6th

Crew 216 Operations Report 06-Dec-19
SOL: 12
Name of person filing report: Mike Lawson
Non-nominal systems: NA
Generator: run
Hours run: 12
From what time last night: 1902
To what time this morning: 0654
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night: 69%
Diesel Reading – 32
Station Propane Reading – 50
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1465756
Water (static tank) (gal): 154
Static to Loft Pump used – yes
Water in Green Hab (gal): 241
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: no
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: yes
Hours: 116.4
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 55
Currently charging: yes
Opportunity rover used: no
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes
Curiosity rover used: yes
Hours: 121.4
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 64%
Currently charging: yes
Notes on rovers: Opportunity and Deimos off-site for maintenance.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:
HabCar used and why, where?
CrewCar used and why, where?
General notes and comments: EVA to Zubrin’s Head. Began station cleaning. Food inventory performed. Samples
returned Robert’s Rock Garden.
Summary of internet: Erratic/unstable wifi connection.
Summary of suits and radios: Suit 4 airflow is weak, possibly a battery charging issue.
Summary of Hab operations: Routine maintenance and housekeeping tasks. Hot water line vibration is extreme when pump
engages, recommend securing it to HAB wall or risk fitting failure.
Summary of GreenHab operations: Routine watering. Greenhab cleaned.
Summary of Science Dome operations: Gypsum experiment completed. Lab cleaned.
Summary of RAM operations: RAM Rover power charging stations inop, electrical issues remain unresolved. RAM cleaned except for
outer airlock. All tools returned.
Summary of any observatory issues: Observatory cleaned.
Summary of health and safety issues: NA
Questions, concerns and requests to Mission Support: 1 bag B, 5 bag NB trash in airlock.

Copyright © The Mars Society. All rights reserved. | Main Site